OpenStudy (anonymous):

Medal and fan! which product represents the solution to the system? 3x+4y=-2 2x+5y=6

OpenStudy (anonymous):

apply matrices. \[\left[\begin{matrix}3 & 4 \\ 2 & 5\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right)=\left(\begin{matrix}-2 \\ 6\end{matrix}\right)\]

OpenStudy (anonymous):

I am sure you can follow the rules of matrices to find the solution of x and y now from the previous post

OpenStudy (anonymous):

my answer^^

OpenStudy (anonymous):

try calculating the inverse of the matrix again

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[\left[\begin{matrix}3 & 4 \\ 2 & 5\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right)=\left(\begin{matrix}-2 \\ 6\end{matrix}\right)\] \[\left(\begin{matrix}x \\ y\end{matrix}\right)=\left[\begin{matrix}3 & 4 \\ 2 & 5\end{matrix}\right]^{-1}\left(\begin{matrix}-2 \\ 6\end{matrix}\right)\]

OpenStudy (anonymous):

remember the inverse of a matrix for a 2x2 matrix \[A=\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] then \[A ^{-1}=\frac{ 1 }{ ad-bc }\left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\]

OpenStudy (anonymous):

feel free to fan me so you can ask me for help straight away when you need it :)